# Invariant Subspace Decomposition

### v1

@@ Abstract
We consider the task of predicting a response $Y$ from a set of covariates $X$ in settings where the conditional distribution of $Y$ given $X$ changes over time. For this to be feasible, assumptions on how the conditional distribution changes over time are required. Existing approaches assume, for example, that changes occur smoothly over time so that short-term prediction using only the recent past becomes feasible. In this work, we propose a novel invariance-based framework for linear conditionals, called Invariant Subspace Decomposition (ISD), that splits the conditional distribution into a time-invariant and a residual time-dependent component. As we show, this decomposition can be utilized both for zero-shot and time-adaptation prediction tasks, that is, settings where either no or a small amount of training data is available at the time points we want to predict $Y$ at, respectively. We propose a practical estimation procedure, which automatically infers the decomposition using tools from approximate joint matrix diagonalization. Furthermore, we provide finite sample guarantees for the proposed estimator and demonstrate empirically that it indeed improves on approaches that do not use the additional invariant structure.
@@ 1 Introduction
Many commonly studied systems evolve with time, giving rise to heterogeneous data. Indeed, changes over time in the data generating process are reflected by shifts in the data distribution, and make it in general impossible to find a fixed model that consistently describes the system through time. In regression settings, varying-coefficients models (e.g., Hastie and Tibshirani, 1993; Fan and Zhang, 2008) are a common approach to deal with dynamic system behaviours related to changes in the functional relationship between a response and some covariates. Another approach for online estimation of changing parameters uses state-space models and filtering solutions (see, for example, Durbin and Koopman, 2012). Changes in time of the covariates distribution can instead be modeled, for example, as non-stationarity in time-series.
In this work, we analyze the problem of estimating the relation between a response $Y_{t}$ and a set of covariates (or predictors) $X_{t}$, $t\in\mathbb{N}$, that are observed over some period of time, with the goal of short term prediction of an unobserved response, when the observations of new covariates become available. Changes in the data distribution can occur, for example, if the mechanism relating the response to the covariates or the covariates distribution vary with time $t\in\mathbb{N}$: in our setting, we allow for both such changes to happen. The task of predicting an unobserved response is then only feasible if some information from past time points can be used. This, in general, requires assumptions on how the underlying data generating model changes over time. For example, if changes happen smoothly, then the most relevant observations are likely the last available. To also exploit information further in the past, it is sometimes of interest to look for invariant patterns that persist through time. In this spirit, Meinshausen and Bühlmann (2015) define the maximin effect as a worst case optimal model that maintains good predictive properties at all observed times. Invariance based methods are also used in the machine learning literature to deal with covariates shifts. One relevant example is unsupervised domain adaptation (see, for example, Sun et al., 2017; Zhao et al., 2019; Stojanov et al., 2021), which aims at transferring invariant information from a source domain to a target domain in which the response is not observed. The concept of invariance is widely explored in the field of causal inference, too. In causal models, changes in the covariates distribution are modeled as interventions, and different interventional settings represent different environments. In this context, Peters et al. (2016) propose to look for a set of stable causal predictors, that is, a set of covariates for which the conditional distribution of the response given such covariates remains invariant across different environments. The same idea is extended by Pfister et al. (2019a) to a sequential setting in which the different environments are implicitly generated by changes of the covariates distribution over time. Such invariances can then be used for prediction in previously unseen environments, too (e.g., Rojas-Carulla et al., 2018; Magliacane et al., 2018; Pfister et al., 2021).
In this work, we build on this invariance idea and exploit it for time adaptation: we propose an invariance based framework, which we call invariant subspace decomposition (ISD), for the estimation of time-varying regression models. More in detail, we aim at finding the time-varying functional relationship between the covariates and the response through the maximization of the explained variance at time $t$, which is given for some function $f$ by $\operatorname{\Delta Var_{t}}(f)\coloneqq\operatorname{Var}(Y_{t})-\operatorname{Var}(Y_{t}-f(X_{t}))$. The explained variance constitutes an objective function equivalent to the mean squared error (with opposite optimization sign), and has the advantage of providing an intuitive evaluation of the predictive quality of a function $f$: negative explained variance indicates in particular that using $f$ for prediction is harmful, in that it is worse than using the zero function. We study the case in which $f$ is a linear function. Key to ISD is the decomposition of the explained variance maximization into a time-invariant and a time-dependent part. To achieve the decomposition, we exploit ideas from independent subspace analysis (see, for example, the work by Gutch and Theis, 2012) and seek a transformation of the original space (in which the covariates $X_{t}$ are observed), such that the number of pairwise uncorrelated components of the transformed covariates is maximized at all observed time points. The key aspect of such a decomposition is that it allows us to look at two sub-problems, each of which has less degrees of freedom than the original one: (i) an invariant problem that can be solved only once exploiting all available heterogeneous data and (ii) an adaptation problem that tunes the invariant solution to a time-point of interest. This result is summarized in Theorem 1. The solution to (i) is a time-invariant linear parameter ($\beta^{\operatorname{inv}}$) that partially describes the data at all times and, in lack of reliable data to solve the time-adaptation step, can be used for zero-shot prediction. The solution to (ii) ($\delta^{\operatorname{res}}_{t}$) can be seen as a fine-tuning of the solution to (i) to learn the time-varying linear parameter of interest ($\gamma_{0,t}$). A simple $2$-dimensional example is shown in Figure 1 and later presented in Example LABEL:ex:running_ex_2d: here, the true time-varying linear parameter varies in a one dimensional subspace of $\mathbb{R}^{2}$, while the time-invariant parameter lies in the corresponding orthogonal subspace. Moreover, the components of the covariates along these two orthogonal subspaces are uncorrelated.
Figure 1: ISD example with covariates of dimension $p=2$. The residual component $\delta^{\operatorname{res}}_{t}$ of the true time-varying parameter $\gamma_{0,t}$—see model (1)—lies on a one-dimensional subspace, visualized by the dots. The invariant component $\beta^{\operatorname{inv}}$ lies instead on the orthogonal subspace. The estimates $\hat{\beta}^{\operatorname{inv}}$ and $\hat{\delta}^{\operatorname{res}}_{t}$ are the ISD solutions to steps (i) and (ii), respectively. Here, $\hat{\beta}^{\operatorname{inv}}$ is estimated using observed time points for $t\in\{-1000,\dots,0\}$ (not shown in the figure). The plot compares the ISD estimates with the OLS solution $\hat{\gamma}_{t}^{\operatorname{OLS}}$ computed on the same data as $\hat{\delta}^{\operatorname{res}}_{t}$, that is, a rolling window of length $16$ (the adaptation data), contained in the time interval $\{1,\dots,350\}$. We provide more details in Example 1.
@@
The fundamental assumption we make for ISD is that the decomposition inferred from available observations generalizes to unseen time points. The main implication of such generalization assumption is that, once an estimate of the invariant component that solves step (i) is available, it can, at all new time points, either be used directly for prediction or as part of a two-step estimation that is fine-tuned with step (ii), which only solves the time adaptation part of the problem.
@@
Figure 2 shows the cumulative explained variance by the invariant component and the adapted invariant component estimated in a simulated setting at training time (left) within the observed time horizon, and at testing time (right) when new observations become available. We compare the ISD with the standard OLS solution, computed both on all training observations and on a rolling window over the latest new observations, and with the maximin effect ($\widehat{\beta}^{\operatorname{mm}}$) by Meinshausen and Bühlmann (2015), which maximizes the worst case explained variance over the available observations.
@@
Figure 2: Example of average explained variance (distribution over 20 runs) obtained at training time (historical data) and of cumulative explained variance obtained testing time (adaptation data) (in one of the 20 runs), where the time-varying components in the historical and adaptation data have disjoint support. The displayed simulation considers predictors of dimension $p=10$ and an invariant component of dimension $7$. As baselines, we use (a) the true time-varying parameter $\gamma_{0,t}$ and (b) the oracle invariant component $\beta^{\operatorname{inv}}$. $n=6000$ historical observations are used to estimate: (c) the invariant component $\hat{\beta}^{\operatorname{inv}}$ of the ISD framework, (d) the OLS solution $\hat{\beta}^{\operatorname{OLS}}$, (e) the maximin effect $\hat{\beta}^{\operatorname{mm}}$. Starting from $t=0$ after the observed history, windows of length $3p$ are used to estimate: (f) the adaptation parameter $\hat{\delta}^{\operatorname{res}}_{t}$ for the oracle parameter $\beta^{\operatorname{inv}}$ (oracle ISD), (g) the adaptation parameter $\hat{\delta}^{\operatorname{res}}_{t}$ for $\hat{\beta}^{\operatorname{inv}}$ (ISD) and (h) the rolling window OLS solution $\hat{\gamma}^{\operatorname{OLS}}_{t}$.
@@
The remainder of the paper is structured as follows. In Section 2 we introduce the ISD framework for orthogonal transformations. We then describe, in Section 3, the two tasks that ISD solves, namely zero-shot prediction and time-adaptation. In Section 4 we propose an estimation method for ISD, and provide finite sample generalization guarantees for the proposed estimator. Finally, in Section 5, we illustrate ISD based on numerical experiments and validate the theoretical results presented in the paper.
@@ 2 Invariant subspace decomposition
Let $(X_{t},Y_{t})_{t\in\mathbb{N}}\subseteq\mathbb{R}^{p}\times\mathbb{R}^{1}$ be a sequence of independent random vectors satisfying for all $t\in\mathbb{N}$ a linear model of the form
$Y_{t}=X_{t}^{\top}\gamma_{0,t}+\epsilon_{t},\quad\text{with }\mathbb{E}[\epsilon_{t}\mid X_{t}]=0,$ (1)
where $\gamma_{0,t}\in\mathbb{R}^{p}$ is the (unknown) true time-varying parameter. We further assume that for all $t\in\mathbb{N}$ the covariance matrix of the predictors $\Sigma_{t}\coloneqq\operatorname{Var}(X_{t})$ is strictly positive definite. Moreover, for all $n\in\mathbb{N}$ we define the set $[n]\coloneqq\{1,\dots,n\}$.
We are interested in predicting the response $Y_{n+m+1}$ from the observed data $(X_{t},Y_{t})_{t\in[n+m]}$ and the predictor $X_{n+m+1}$ in settings where the true time-varying parameter is constant or approximately constant only in short periods of time, e.g., step-wise constant (we refer to the case $m=0$ as zero-shot generalization; $m>0$ is a special case of what we refer to as time-adaptation, see Section 3). A naive solution is to only consider the most recent observations, assuming that the model is approximately time-invariant over these observations. Our goal in this work is to improve on this approach and exploit all observations in $[n]$. To achieve this, we propose to learn two parameters whose sum equals the time-varying parameter $\gamma_{0,t}$: a time-invariant parameter $\beta^{\operatorname{inv}}\in\mathbb{R}^{p}$, defined formally below in Section 2.2, and an adaptation parameter $\delta^{\operatorname{res}}_{t}$ defined formally in Section 2.3 that explains the remaining residuals $Y_{t}-X_{t}^{\top}\beta$. The separation allows us to first use all available observations in $[n]$ to estimate the time-invariant parameter, and to then improve the prediction (using only the most recent observations) by learning an adaptation parameter with less degrees of freedom, explaining the remaining variance.
Before we define time-invariant parameters, we first consider the explained variance for a parameter $\beta\in\mathbb{R}^{p}$ at time $t\in\mathbb{N}$, defined by
@@
The true time-varying parameter can always be expressed as
@@
We show in Theorem 1 that, under some additional assumptions, the true time-varying parameter $\gamma_{0,t}$ can be decomposed into a time-invariant component $\beta^{\operatorname{inv}}$ and a time-varying component $\delta^{\operatorname{res}}$; furthermore, this decomposition can be obtained by splitting the parameter space into two orthogonal linear subspaces $\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\subseteq\mathbb{R}^{p}$ with $\mathcal{S}^{\operatorname{inv}}\oplus\mathcal{S}^{\operatorname{res}}=\mathbb{R}^{p}$ such that, for all $t\in\mathbb{N}$,
[…]
In this work, we consider explained variance as an objective. Equivalently, one can consider the mean squared prediction error (MSPE).

### v2

@@ Abstract
We consider the task of predicting a response $Y$ from a set of covariates $X$ in settings where the conditional distribution of $Y$ given $X$ changes over time. For this to be feasible, assumptions on how the conditional distribution changes over time are required. Existing approaches assume, for example, that changes occur smoothly over time so that short-term prediction using only the recent past becomes feasible. To additionally exploit observations further in the past, we propose a novel invariance-based framework for linear conditionals, called Invariant Subspace Decomposition (ISD), that splits the conditional distribution into a time-invariant and a residual time-dependent component. As we show, this decomposition can be utilized both for zero-shot and time-adaptation prediction tasks, that is, settings where either no or a small amount of training data is available at the time points we want to predict $Y$ at, respectively. We propose a practical estimation procedure, which automatically infers the decomposition using tools from approximate joint matrix diagonalization. Furthermore, we provide finite sample guarantees for the proposed estimator and demonstrate empirically that it indeed improves on approaches that do not use the additional invariant structure.
Keywords: invariance, distribution shift, time adaptation, zero-shot learning, joint block diagonalization.
@@ 1 Introduction
Many commonly studied systems evolve with time, giving rise to heterogeneous data. Indeed, changes over time in the data generating process lead to shifts in the observed data distribution, and make it in general impossible to find a fixed model that consistently describes the system over time.
In this work, we analyze the problem of estimating the relation between a response $Y_{t}$ and a set of covariates (or predictors) $X_{t}$, $t\in\mathbb{N}$, both of which are observed over some period of time, with the goal of predicting an unobserved response, when new observations of the covariates become available. Two types of distribution shifts across time that can arise are (i) variations in the mechanism relating the response to the covariates and (ii) variations in the covariate distribution. Under distribution shifts, successfully predicting an unobserved response from covariates requires assumptions on how the underlying data generating model changes over time.
In regression settings, varying-coefficients models (e.g., Hastie and Tibshirani, 1993; Fan and Zhang, 2008) are a common approach to deal with dynamic system behaviours related to changes in the functional relationship between a response and some covariates. In these works, the model coefficients are usually assumed to change smoothly, and smoothing methods are used in their estimation. Another approach for online estimation of smoothly changing parameters uses state-space models and filtering solutions (see, for example, Durbin and Koopman, 2012). In general, smooth changes or other kinds of structured variations such as, for example, step-wise changes, allow us to learn a regression function using previous time points from the recent past. To exploit information further in the past, one can look for invariant patterns that persist through time. In this spirit, Meinshausen and Bühlmann (2015) define the maximin effect as a worst case optimal model that maintains good predictive properties at all observed times.
Distribution shifts in time-series models are conceptually similar to distribution shifts across domains. In both cases a key problem is to find invariant parts of the observed data distribution and transfer it to the new domain or time-point. Unsupervised domain adaptation methods aim to predict an unobserved response in the target domain: for the problem to be tractable, they rely, for example, on the invariance of the conditional distribution of the response given the covariates (Sun et al., 2017; Zhao et al., 2019) or on independently changing factors of the joint distribution of the covariates and the response (Zhang et al., 2015; Stojanov et al., 2021). Solutions to the unsupervised domain adaptation problem are often based on aligning the source and target domains by minimizing the discrepancy either between the covariates’ distributions (Zhang et al., 2015) or between the covariates’ second order moments (Sun et al., 2017), or by finding low-dimensional invariant transformations of the variables whose distributions match in the domains of interest (Zhao et al., 2019; Stojanov et al., 2021). Other unsupervised domain adaptation approaches such as the one proposed by Bousmalis et al. (2016) rely on learning disentangled representations where disentangled mean that changes in one representation do not affect the remaining ones. This approach is also explored in non-stationary settings, when changes in the covariate distribution happen over time instead of across different domains, for example in the works by Yao et al. (2022); Li et al. (2024). The goal of these works is not prediction, but rather the identification of generalizable latent states that could be used for arbitrary downstream tasks.
The field of causality widely explores the concept of invariant and independently changing mechanisms, too. In causal models, changes in the covariate distribution are modeled as interventions, and different interventional settings represent different environments. In this context, Peters et al. (2016) propose to look for a set of stable causal predictors, that is, a set of covariates for which the conditional distribution of the response given such covariates remains invariant across different environments. The same idea is extended by Pfister et al. (2019a) to a sequential setting in which the different environments are implicitly generated by changes of the covariate distribution over time. Such invariances can then be used for prediction in previously unseen environments, too (e.g., Rojas-Carulla et al., 2018; Magliacane et al., 2018; Pfister et al., 2021). Other works on causal discovery from heterogeneous or nonstationary data, such as the ones by Huang et al. (2020); Günther et al. (2023), study settings in which the skeleton of the causal graph remains invariant through different contexts or time points, but the causal mechanisms are also allowed to change. Their goal is then to identify the causal graph on observed covariates. To detect whether a mechanism changes, the context or time is included as an additional variable in the graph. As the context or time can be assumed to be exogenous, this often leads to additional identification of causal edges in the graphs. Feng et al. (2022) develop and apply similar ideas in a reinforcement learning setting. Their interest is in learning the causal graph and improving the efficiency of non-stationarity adaptation by disentangling the system variations as latent parameters.
In this work, we build on invariance ideas both from the maximin framework proposed by Meinshausen and Bühlmann (2015) and from the covariates shifts literature, and exploit them for time adaptation. We propose an invariance-based framework, which we call invariant subspace decomposition (ISD), to estimate time-varying regression models in which both the covariate distribution and their relationship with a response can change. In particular, we consider a sequence of independent random vectors $(X_{t},Y_{t})_{t\in\mathbb{N}}\subseteq\mathbb{R}^{p}\times\mathbb{R}$ satisfying for all time points $t\in\mathbb{N}$ a linear model of the form
$Y_{t}=X_{t}^{\top}\gamma_{0,t}+\epsilon_{t}$ (1)
with $\mathbb{E}[\epsilon_{t}\mid X_{t}]=0$, where $\gamma_{0,t}$ is the (unknown) true time-varying parameter. We allow the covariance matrices $\operatorname{Var}(X_{t})$ to change over time, assuming that they are approximately constant in small time windows. Regarding changes in $\gamma_{0,t}$, we only need to assume smoothness during test time (if there is more than one test point)—we provide more details later in Remark 2. Having observed the $X_{t}$ and $Y_{t}$ up until some time point $n$, our goal is to learn $\gamma_{0,t^{*}}$ for some $t^{*}>n$, which we then use to predict $Y_{t^{*}}$ from $X_{t^{*}}$. We propose to do so by considering the explained variance, which, for all $t$, is given for some function $f$ by $\operatorname{\Delta Var}_{t}(f)\coloneqq\operatorname{Var}(Y_{t})-\operatorname{Var}(Y_{t}-f(X_{t}))$. Using the explained variance as the objective function provides an intuitive evaluation of the predictive quality of a function $f$: negative explained variance indicates in particular that using $f$ for prediction is harmful, in that it is worse than the best constant function. Under model (1), $f$ is a linear function fully characterized by a parameter $\beta\in\mathbb{R}^{p}$, i.e., $f(X_{t})=X_{t}^{\top}\beta$, and the true time-varying parameter can always be expressed as
Key to ISD is the decomposition of the explained variance maximization into a time-invariant and a time-dependent part: we show in Theorem 1 that, under some assumptions, for all $t\in\mathbb{N}$, $\gamma_{0,t}$ can be expressed as
where $\overline{\operatorname{\Delta Var}}(\beta)\coloneqq\frac{1}{n}\sum_{t=1}^{n}\operatorname{\Delta Var}_{t}(\beta)$, and $\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\subseteq\mathbb{R}^{p}$ are two orthogonal linear subspaces, with $\mathcal{S}^{\operatorname{inv}}\oplus\mathcal{S}^{\operatorname{res}}=\mathbb{R}^{p}$, assumed to be uniquely determined by the distribution of $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$. The decomposition of $\mathbb{R}^{p}$ into $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ from observed data is achieved by exploiting ideas from independent subspace analysis (e.g., Gutch and Theis, 2012). Unlike works on latent representations, we do not reduce the overall dimensionality of the problem but partition the observed space into two lower dimensional subspaces, $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$. This in particular implies that the two sub-problems in (3) each have less degrees of freedom than the original one. The first sub-problem can be solved using all available heterogeneous observations, which we call historical data: its solution $\beta^{\operatorname{inv}}$ is a time-invariant linear parameter that partially describes the dependence of $Y$ on $X$ at all times. The second sub-problem is a time adaptation problem that tunes the invariant component to a time point of interest: its solution $\delta^{\operatorname{res}}_{t}$ explains the residuals $Y_{t}-X_{t}^{\top}\beta^{\operatorname{inv}}$; the sum $\beta^{\operatorname{inv}}+\delta_{t}^{\operatorname{res}}$ gives an estimator for the time-varying linear parameter of interest $\gamma_{0,t}$. In order to estimate the residual component $\delta_{t}^{\operatorname{res}}$, we assume the model is approximately stationary in a small time window preceding $t$ and use this local subset of the data, which we call adaption data, for estimation. We distinguish two tasks: (i) the zero-shot task, where adaption data is not available and we approximate $\gamma_{0,t}$ by $\beta^{\operatorname{inv}}$ and (ii) the time-adaption task, where adaption data is available and we solve both sub-problems to approximate $\gamma_{0,t}$ by $\beta^{\operatorname{inv}}+\delta_{t}^{\operatorname{res}}$. A two-dimensional example of $\gamma_{0,t}$ and its ISD estimates is shown in Figure 1. The same example is presented again in more detail in Example LABEL:ex:running_ex_2d below and used as a running example throughout the paper.
Figure 1: Example of two-dimensional true parameter $\gamma_{0,t}$ varying on a one-dimensional subspace of $\mathbb{R}^{2}$, and its estimates using ISD (right) compared to rolling window OLS (left). Time is visually encoded using a color map. (Left) True parameter $\gamma_{0,t}$ (hexagons) on $350$ test points and OLS estimates $\hat{\gamma}^{\operatorname{OLS}}_{t}$ based on rolling windows of size $16$. (Right) Same test data and true parameters $\gamma_{0,t}$, but now we additionally use $1000$ prior time-points as historical data (not shown) to estimate the decomposition of $\mathbb{R}^{2}$ into the orthogonal subspaces $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ (dashed lines). Next, we estimate $\beta^{\operatorname{inv}}$ using the historical data and $\hat{\mathcal{S}}^{\operatorname{inv}}$. Then, using the same rolling windows as in the left plot as adaption data, we estimate $\delta^{\operatorname{res}}_{t}$ using $\hat{\mathcal{S}}^{\operatorname{res}}$. The ISD estimates are then given by $\hat{\gamma}_{t}^{ISD}=\hat{\beta}^{\operatorname{inv}}+\hat{\delta}^{\operatorname{res}}_{t}$. All details on the generative model are provided in Example 1. The subspaces $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ do not need to be axis aligned, so ISD is applicable even in cases where the conditional of $Y_{t}$ given $X_{t}$ and all conditionals of $Y_{t}$ given subsets of $X_{t}$ vary over time.
@@
The fundamental assumption we make to apply the ISD framework to new data (Assumption 2) is that the decomposition inferred from available observations generalizes to unseen time points. This guarantees that the estimated invariant component can be meaningfully used for prediction at all new time points, either directly or as part of a two-step estimation that is fine-tuned by solving the time adaptation sub-problem.
@@
The ISD framework allows us to improve on naive methods that directly maximize the explained variance on $\mathbb{R}^{p}$ using only the most recent available observations and on methods focusing on invariance across environments or time points such as the maximin by Meinshausen and Bühlmann (2015), which do not account for time-adaptation. We show in particular in Theorem 3 that isolating an invariant component and reducing the dimensionality of the adaptation problem guarantees lower prediction error or, equivalently, higher explained variance compared to naive methods. An example is provided for a simulated setting in Figure 2, which shows on the left the average explained variance by the invariant component computed at training time on historical data, and on the right the cumulative explained variance by the invariant (zero-shot task) and the adapted invariant component (time-adaptation task) at testing time when new observations become available. For the zero-shot task, we compare the ISD invariant component with the standard OLS solution, computed on all training observations, and with the maximin effect ($\hat{\beta}^{\operatorname{mm}}$) by Meinshausen and Bühlmann (2015), which maximizes the worst case explained variance over the available observations. For the time-adaptation task, we compare the ISD with the standard OLS solution computed on a rolling window over the latest new observations. For both tasks, the ISD estimates achieve higher cumulative explained variance at new time-points.
@@
Figure 2: For the data-generating model in Section 5.2, we plot (left) the average explained variance (distribution over 20 runs) obtained at training time (historical data) and (right) the cumulative explained variance obtained testing time (adaptation data) (in one of the 20 runs); in this example, the time-varying components in the historical and adaptation data have disjoint support. The example considers $p=10$-dimensional predictors and an invariant component of dimension $7$. As baselines, we use (i) the true time-varying parameter $\gamma_{0,t}$, which maximizes the explained variance at all observed time points $t$, and (ii) the oracle invariant component $\beta^{\operatorname{inv}}$. $6000$ historical observations are used to estimate: (iii) the invariant component $\hat{\beta}^{\operatorname{inv}}$ of the ISD framework, (iv) the OLS solution $\hat{\beta}^{\operatorname{OLS}}$, (v) the maximin effect $\hat{\beta}^{\operatorname{mm}}$. Starting from $t=0$ after the observed history, windows of length $3p$ are used to estimate: (vi) the adaptation parameter $\hat{\delta}^{\operatorname{res}}_{t}$ for $\hat{\beta}^{\operatorname{inv}}$ to obtain the ISD estimate $\hat{\gamma}^{ISD}$ and (vii) the rolling window OLS solution $\hat{\gamma}^{\operatorname{OLS}}_{t}$. While at training time on historical data the ISD invariant component $\hat{\beta}^{\operatorname{inv}}$ is the most conservative, with the lowest average explained variance, after a distribution shift (adaptation data) the same component can explain higher variance than other methods based on historical data only ($\hat{\beta}^{\operatorname{OLS}},\hat{\beta}^{\operatorname{mm}}$), and can be tuned to new time points to improve on estimators based on adaptation data only ($\hat{\gamma}^{\operatorname{OLS}}_{t}$).
@@
The remainder of the paper is structured as follows. In Section 2 we introduce the ISD framework, starting from the identification of the invariant and residual subspaces for orthogonal covariates transformation and showing the explained variance separation (3) in Theorem 1. In Sections 2.2 and 2.3 we define population estimators for the invariant and residual components, solutions to the two sub-problems in (3). We then describe, in Section 3, the two tasks that ISD solves, namely zero-shot prediction and time-adaptation, and provide a characterization of the invariant component as a worst-case optimal parameter. In Section 4 we propose an estimation method for ISD, and provide finite sample generalization guarantees for the proposed estimator. Finally, in Section 5, we illustrate ISD based on numerical experiments, both on simulated and on real world data, and validate the theoretical results presented in the paper.
@@ 2 Invariant subspace decomposition
We formalize the setup described in the introduction in the following setting.
Setting 1.
Let $(X_{t},Y_{t})_{t\in\mathbb{N}}\subseteq\mathbb{R}^{p}\times\mathbb{R}^{1}$ be a sequence of independent random vectors satisfying for all $t\in\mathbb{N}$ a linear model as in (1), with $\gamma_{0,t}\in\mathbb{R}^{p}$ and $\mathbb{E}[\epsilon_{t}\mid X_{t}]=0$. Assume that for all $t\in\mathbb{N}$ the covariance matrix of the predictors $\Sigma_{t}\coloneqq\operatorname{Var}(X_{t})$ is strictly positive definite. Moreover, for all $n\in\mathbb{N}$, let $[n]\coloneqq\{1,\dots,n\}$ and assume we observe $n$ observations $(X_{t},Y_{t})_{t\in[n]}$ from model (1), which we call historical data. Additionally, let $\mathcal{I}^{\operatorname{ad}}\subseteq\mathbb{N}$ be an interval of consecutive time points with $m:=|\mathcal{I}^{\operatorname{ad}}|>p$ and $\min_{t\in\mathcal{I}^{\operatorname{ad}}}t>n$. Assume we observe a second set of observations $(X_{t},Y_{t})_{t\in\mathcal{I}^{\operatorname{ad}}}$ from the same model but succeeding the historical data, which we call adaptation data. Finally, denote by $t^{*}>\max_{t\in\mathcal{I}^{\operatorname{ad}}}$ a time point of interest that occurs after the adaptation data and assume that, for all $t\in\mathcal{I}^{\operatorname{ad}}\cup\{t^{*}\}$, the quantities $\gamma_{0,t}$, $\Sigma_{t}$and $\operatorname{Var}(\epsilon_{t})$ in model (1) are constant (in practice, being approximately constant is sufficient).
Our goal is to predict $Y_{t^{*}}$ from $X_{t^{*}}$. A naive solution is to only consider the adaptation data in $\mathcal{I}^{\operatorname{ad}}$ (see also the analysis in Section 3): we aim to improve on this approach by additionally exploiting historical data in $[n]$. In particular, depending on whether we only have access to the historical data or we additionally have access to the data in $\mathcal{I}^{\operatorname{ad}}$, we can solve (i) the zero-shot task or (ii) the adaptation task, respectively. The full setup including the different tasks is visualized in Figure 3.
[…]
The two tasks correspond to the two sub-problems in (3): the solution to (i) is a time-invariant parameter $\beta^{\operatorname{inv}}\in\mathbb{R}^{p}$, defined formally below in Section 2.2; the solution to (ii) is an adaptation parameter $\delta^{\operatorname{res}}_{t}$ formally defined in Section 2.3 that satisfies $\beta^{\operatorname{inv}}+\delta^{\operatorname{res}}_{t}=\gamma_{0,t}$.
We start by defining what a time-invariant parameter is. To do so, we first consider the explained variance for a parameter $\beta\in\mathbb{R}^{p}$ at time $t\in\mathbb{N}$, defined by
@@
We use this definition, in Section 2.1, to define the subspaces $\mathcal{S}^{\operatorname{inv}}$ and $\mathcal{S}^{\operatorname{res}}$ that allow us to obtain the separation of $\gamma_{0,t}$ as in (3). In particular, as we show in Proposition 2, the definition of the invariant subspace $\mathcal{S}^{\operatorname{inv}}$ guarantees that the invariant component $\beta^{\operatorname{inv}}$ maximizing the pooled explained variance over $\mathcal{S}^{\operatorname{inv}}$ is always a time-invariant parameter according to Definition 1.
While we consider the explained variance here, one can equivalently consider the mean squared error (MSPE).